# Purely gravitational dark matter production in warm inflation

### v1

@@ Abstract
We consider an appealing scenario for the production of purely gravitational dark matter in the background of warm inflation, a mechanism that maintains stable thermal bath during inflation. Through systematic investigation of various gravitational production channels, we reveal distinctive features compared to the standard inflation scenario. Notably, the inflaton annihilation channel in warm inflation exhibits markedly different thermodynamics from the standard inflation paradigm, leading to a suppression on the production of sub-inflaton-mass dark matter. For the production channel of inflationary vacuum fluctuations, we find a correlation of $\rho_{\chi}\propto m_{\chi}^{5/2}$ for the conformally coupled dark matter, which expands the feasible range of dark matter mass. Our results also indicates that a minimum temperature threshold of $10^{-6}M_{P}$ is necessary for warm inflation, which allows adequate dark matter production. With observational constraints, our results provide stringent limits on the mass range of purely gravitational dark matter with sufficient density: $10^{-8}-10^{-2}M_{P}$ for minimal coupling and $10^{-14}-10^{-2}M_{P}$ for conformal coupling.
@@ I Introduction
Production of purely gravitational DM generally occurs in the very early universe, e.g., during inflation and following reheating stage. There are mainly two feasible production mechanisms. One is called the cosmological gravitational particle production (CGPP) [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 12, 13, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52], which refers to the particle production from the vacuum of an expanding universe. This mechanism arises from the temporal variation in the vacuum state of the quantum field in the cosmological background. The vacuum state at the previous moment can evolve into an excited state relative to the vacuum state at the next moment, indicating the particle production from the vacuum. There is a detailed review article discussed the CGPP of DM [23]. Another mechanism is to connect the thermal bath and DM through graviton as a medium, which is also known as the gravitational DM freeze-in [53, 54, 55, 17, 56, 16, 57, 14, 15]. Efficiency of this mechanism strongly depends on the temperature of the thermal bath. In the scenario of reheating, it necessitates an exceptionally efficient reheating mechanism to achieve a sufficiently high temperature for DM freeze-in.
Reheating is a conventional scenario that connects the inflationary phase and the radiation-dominant (RD) phase in the standard inflation (SI) theory [58, 59, 60, 61, 62, 63]. However, it is not the only possibility. An alternative scenario suggests that the inflaton field may have a strong interaction with radiation during inflation, thereby dissipating energy into radiation to sustain a stable thermal bath, up until the universe completes the transition from inflation to the RD stage. It is known as the warm inflation (WI) paradigm [64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85]. There are many motivations for WI. For instance, prior studies indicate that WI with strong dissipation can inhibit super-Planckian field variation for the inflaton [86], thereby facilitating the realization of small-field inflation theory . This is a nice feature from the perspective of effective field theory. Furthermore, WI is capable of producing an enhanced scalar perturbation spectrum, which consequently suppresses the tensor-to-scalar ratio of the primordial perturbations [74, 75]. As a result, it revives many inflationary models previously deemed unfavorable, such as $\phi^{2}$ and $\phi^{4}$ potentials, which have been excluded in the SI paradigm by Planck and BICEP/$Keck$ results [87, 88].
@@
We are tremendously curious about the gravitational freeze-in mechanism of DM in the WI background. Considering the gravitational annihilation of SM particles and inflatons, this study demonstrates the dependence of DM yield on the WI temperature and DM mass, indicating a suppression on the production of sub-inflaton-mass DM. The contribution of CGPP mechanism is also taken into account, revealing different yield-mass correlations from SI. In addition, the combined observational constraints to the parameter space will also be presented, showing stringent limits on the mass range of DM.
@@ II Background dynamics of warm inflation
into Eqs. (6), (7), and (8), the background dynamics of WI can be numerically solved. Let us now choose several sets of parameters to demonstrate the evolution of various physical quantities during the late stage of WI. The results are depicted in Fig. 1. Here we fix $m_{\phi}=10^{-5}M_{P}$, $p=0$, and set $c=1(-1)$ for the left(right) panel of the figure. The solid and dashed curves signify different values of $C_{\Upsilon}$, or different initial value of $Q$, respectively. It is evident that the radiation sustains an almost constant temperature with $T\gg H$ until the end of inflation. This is an important characteristic of WI that will be used in this study. Note that in the WI scenario, the end of inflation is driven by the energy density of radiation $\rho_{r}$ surpassing that of the inflaton $\rho_{\phi}$, rendering the $\epsilon_{V}$ criterion used in the SI framework inapplicable. Only the $\epsilon_{H}$ criterion is valid in WI, that is, WI ends when $\epsilon_{H}\equiv-\dot{H}/H^{2}>1$, or equivalently, $\ddot{a}<0$.
Figure 1: Evolution of various dimensionless quantities during the late stage of WI. The inflationary potential is assumed to possess a $\phi^{2}$ bottom with $m_{\phi}=10^{-5}M_{P}$. In the left panel we set $p=0$, $c=1$, and $C_{\Upsilon}=0.1(10)$ for the solid(dashed) curves. In the right panel we set $p=0$, $c=-1$, and $C_{\Upsilon}=2\times 10^{-8}(2\times 10^{-7})$ for the solid(dashed) curves. The end of WI is fixed at $N=0$.
@@ III Gravitational dark matter production
$+\frac{1}{M_{P}}\left[h_{\mu\nu}T^{\mu\nu}_{M}+\mathcal{L}_{3}(h)\right]+...\Big{\}}.$ (14)
Here, ~$\tilde{\mathcal{L}}_{M}$ refers to the material fields in the FLRW background, $\mathcal{L}_{n}(h)$ refers to the $n$-th order term of $h_{\mu\nu}$, and $T^{\mu\nu}_{M}$ is the energy-momentum tensor of material fields. If we consider the perturbative quantization of gravity, then $h_{\mu\nu}$ can be regarded as graviton, and $h_{\mu\nu}T^{\mu\nu}_{M}$ describes the leading order interaction between graviton and matter.
The action Eq. (III) actually divides DM production into two channels. One is the CGPP mechanism mentioned earlier, which is induced by the DM part in ~$\tilde{\mathcal{L}}_{M}$. The other is the freeze-in mechanism through the graviton-mediated interactions between DM and other particles, induced by $h_{\mu\nu}T^{\mu\nu}_{M}$ terms. The interaction here is mainly the $s$-channel annihilation process, as shown in Fig. 2. During WI, both radiation particles and inflatons possess considerable energy due to the high temperature, which significantly enhances the gravitational scattering cross-section, thereby facilitating the production of massive DM particles.
@@ III.1 Cross-section of graviton-mediated annihilation
here the factor of $1/2$ is considered for neutrino since they are Weyl particles.
@@ IV.2 Annihilation of inflatons
In contrast to the annihilation of inflatons in the SI case with the same $m_{\phi}$, the WI case exhibit two significant differences. Firstly, the DM with $m_{\chi}>m_{\phi}$ can be produced in the WI case due to the quasi-thermal state of inflatons, while it is prohibited in the SI case. Secondly, if $\phi_{e}\sim 0.1-1M_{P}$ is estimated at the end of WI, we will find that the yield of DM with $m_{\chi}\ll m_{\phi}$ in the WI case is roughly $10^{-6}$ lower than that in the SI case with $T_{\mathrm{reh}}=T_{e}$. It reveals the enormously higher annihilation efficiency of static inflatons than moving inflatons. To understand this, we take $\rho_{r}\simeq\rho_{\phi}$ at the end of WI and rewrite Eq. (43) as:
@@
Then comparing it with Eq. (41), we observe that the interaction term for WI is suppressed by a factor of $1/g_{*}^{2}\sim 10^{-4}$. Moreover, as we mentioned earlier, the WI models typically have a reduced $\phi$ variation, which means $\phi_{e}^{\mathrm{WI}}<\phi_{e}^{\mathrm{SI}}$ or $\rho_{\phi}^{\mathrm{WI}}<\rho_{\phi}^{\mathrm{SI}}$ at the end of inflation. Therefore, the interaction term is further suppressed by several orders of magnitude, so does the DM yield.
@@ IV.3 Contribution of CGPP mechanism
We adopt a $\phi^{2}$ WI model with mass parameters $m_{\phi}=10^{-5}M_{P}$ and $10^{-7}M_{P}$. By adjusting the parameters in Eq. (10), we obtain WI backgrounds with $T_{e}=5\times 10^{-4}M_{P}$, $H_{e}=10^{-6}M_{P}$ and $T_{e}=5\times 10^{-5}M_{P}$, $H_{e}=10^{-8}M_{P}$, respectively. The CGPP of DM with minimal coupling ($\xi=0$) and conformal coupling ($\xi=1/6$) is numerically investigated. The results are shown in Fig. 5. These results are actually indicative of the general characteristics of a wide range of WI models, although they are based on a specific model.
@@
Here we present a threshold for $T_{e}$. If WI takes place with $T_{e}$ below this threshold, there will not be an overproduction of DM for any mass. By comparison, the SI scenario also has a correlation of $\rho_{\chi,0}\propto m_{\chi}^{1/2}$ for minimal coupling. However, for conformal coupling, the situation is not exactly the same. Due to the existence of reheating process, there are different correlations for two ranges of $m_{\chi}$ in the SI scenario:
@@
where $H_{\mathrm{reh}}$ is the Hubble parameter when reheating ends. To a certain extent, the WI is similar to the SI with instantaneous reheating, corresponding to $H_{\mathrm{reh}}=H_{e}$. Therefore, in WI, the $\rho_{\chi,0}\propto m_{\chi}^{2}$ interval is nonexistent for conformal coupling. This actually narrows the DM mass range of overproduction compared with the SI scenario when other conditions keep the same, thereby expanding the feasible DM mass range.
Figure 5: Relic abundance of scalar DM with minimal (red curves) and conformal (blue curves) coupling via CGPP mechanism. The inflationary background is served by a $\phi^{2}$ WI model. We set $m_{\phi}=10^{-5}(10^{-7})M_{P}$, $T_{e}=5\times 10^{-4}(5\times 10^{-5})M_{P}$, and $H_{e}=10^{-6}(10^{-8})M_{P}$ for the solid(dashed) curves.
For the $m_{\chi}>H_{e}$ case, apparently, a little increase in mass leads to a rapid decrease in the relic abundance. The super-Hubble-mass DM with $m_{\chi}\sim 10-30H_{e}$ can be sufficiently produced and constitutes the entirety of observed DM abundance. This is similar to the SI case.
@@ V Observational constraints
where $N_{*}\sim 60$ denotes the $e$-folding number between the $k_{*}$ crossing and the end of inflation. $N_{\mathrm{tot}}$ denotes the total $e$-folding number during the inflationary phase, and its value is dependent on the choice of the infrared cut-off of $k$. Consistent with previous discussions, we set the cut-off as $k>a_{0}H_{0}$, which leads to $N_{\mathrm{tot}}\sim 65$. Combining the observational constraint and the above equations, we restrict that $\alpha>0.089$. For scalar DM with conformal coupling, this condition is inherently fulfilled, thereby eliminating any constraints on its mass. However, for the minimally coupled scalar DM, this condition implies that $m_{\chi}>0.52H_{*}\sim m_{\phi}$, which implies that only DM with a super-Hubble scale mass is allowed in this case. This constraint is depicted by the red dashed line in Fig. 6. The mass ranges of DM with sufficient density are limited to $10^{-8}-10^{-2}M_{P}$ for minimal coupling and $10^{-14}-10^{-2}M_{P}$ for conformal coupling.
@@ VI Conclusion
We have investigated the purely gravitational DM production within the WI background. Three possible gravitational channels for DM production are explored, and exhibit distinctive features compared to the SI scenario. For the freeze-in channel via the SM particles annihilation, we derive the interaction term in the Boltzmann equation and calculate the DM yield. The analytical approach is utilized for the light DM case, which is consistent with the numerical results. For the inflaton annihilation channel, we obtain markedly different results from the SI scenario. Due to the quasi-thermal state of inflaton, the channel of producing super-inflaton-mass DM is open in WI, while the sub-inflaton-mass case is suppressed compared to the SI scenario. In the context of the CGPP channel, we identify the temperature threshold for sufficient DM production. We also show a correlation of $\rho_{\chi,0}\propto m_{\chi}^{5/2}$ for the conformally coupled DM, indicating differences from the SI scenario. Furthermore, the constraints on the DM mass $m_{\chi}$ and the temperature $T_{e}$ at the end of WI is explored. We demonstrate the feasible parameter space limited by the observations of DM abundance and isocurvature perturbations. The results of this study can be extended to purely gravitational fermion and vector DM, as they are similar to the scalar DM with conformal and minimal coupling, respectively.

### v3

@@ Abstract
We consider an appealing scenario for the production of purely gravitational dark matter in the background of warm inflation, a mechanism that maintains stable thermal bath during inflation. Through systematic investigation of various gravitational production channels, we reveal distinctive features compared to the standard inflation scenario. Notably, the inflaton annihilation channel in warm inflation exhibits markedly different thermodynamics from the standard inflation paradigm, leading to a suppression on the production of sub-inflaton-mass dark matter. For the production channel of inflationary vacuum fluctuations, we find an abundance-mass correlation of $\rho_{\chi}\propto m_{\chi}^{1/2}(m_{\chi}^{5/2})$ for the sub-Hubble-mass dark matter with minimal(conformal) coupling. Our results also indicate that a minimum temperature threshold of $10^{-6}M_{P}$ is necessary for warm inflation, which allows adequate dark matter production. With observational constraints, our results provide stringent limits on the mass range of purely gravitational dark matter with sufficient density: $10^{-8}-10^{-2}M_{P}$ for minimal coupling and $10^{-14}-10^{-2}M_{P}$ for conformal coupling.
@@ I Introduction
Production of purely gravitational DM generally occurs in the very early universe, e.g., during inflation and following reheating stage. Two feasible production mechanisms are usually discussed. One is called the cosmological gravitational particle production (CGPP) [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 12, 13, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52], which refers to the particle production from the vacuum of an expanding universe. This mechanism arises from the temporal variation in the vacuum state of the quantum field in the cosmological background. The vacuum state at the previous moment can evolve into an excited state relative to the vacuum state at the next moment, indicating the particle production from the vacuum. There is a detailed review article discussed the CGPP of DM [23]. Another mechanism is to connect the thermal bath and DM through graviton as a medium, which is also known as the gravitational DM freeze-in [53, 54, 55, 17, 56, 16, 57, 14, 15, 58]. Efficiency of this mechanism strongly depends on the temperature of the thermal bath. In the scenario of reheating, it necessitates an exceptionally efficient reheating mechanism to achieve a sufficiently high temperature for DM freeze-in.
Reheating is a conventional scenario that connects the inflationary phase and the radiation-dominant (RD) phase in the standard inflation (SI) theory [59, 60, 61, 62, 63, 64]. However, it is not the only possibility. An alternative scenario suggests that the inflaton field may have a strong interaction with radiation during inflation, thereby dissipating energy into radiation to sustain a stable thermal bath, up until the universe completes the transition from inflation to the RD stage. It is known as the warm inflation (WI) paradigm [65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91]. There are many motivations for WI. For instance, prior studies indicate that WI with strong dissipation can inhibit super-Planckian field variation for the inflaton [92], thereby facilitating the realization of small-field inflation theory. This is a nice feature from the perspective of effective field theory. Furthermore, WI is capable of producing an enhanced scalar perturbation spectrum, which consequently suppresses the tensor-to-scalar ratio of the primordial perturbations [75, 76]. As a result, it revives many inflationary models previously deemed unfavorable, such as $\phi^{2}$ and $\phi^{4}$ potentials, which have been excluded in the SI paradigm by Planck and BICEP/$Keck$ results [93, 94].
@@
We are curious about the gravitational freeze-in mechanism of DM in the WI background. Considering the gravitational annihilation of SM particles and inflatons, this study demonstrates the dependence of DM yield on the WI temperature and DM mass, indicating a suppression on the production of sub-inflaton-mass DM. The contribution of CGPP mechanism is also taken into account, revealing different abundance-mass correlations from SI. In addition, the combined observational constraints to the parameter space will also be presented, showing stringent limits on the mass range of DM.
@@ II Background dynamics of warm inflation
into Eqs. (6), (7), and (8), the background dynamics of WI can be numerically solved. Let us now choose several sets of parameters to demonstrate the evolution of various physical quantities during the late stage of WI. The results are depicted in Fig. 1. Here we fix $m_{\phi}=10^{-5}M_{P}$, $p=0$, and set c=1(1)c=1(-1) for the left(right) panel of the figure. The solid and dashed curves signify different values of $C_{\Upsilon}$, or different initial value of $Q$, respectively. It is evident that the radiation sustains an almost constant temperature with $T\gg H$ until the end of inflation. This is an important characteristic of WI that will be used in this study. Note that in the WI scenario, the end of inflation is driven by the energy density of radiation $\rho_{r}$ surpassing that of the inflaton $\rho_{\phi}$, rendering the $\epsilon_{V}$ criterion used in the SI framework inapplicable. Only the $\epsilon_{H}$ criterion is valid in WI, that is, WI ends when $\epsilon_{H}\equiv-\dot{H}/H^{2}>1$, or equivalently, $\ddot{a}<0$.
Figure 1: Evolution of various dimensionless quantities during the late stage of WI. The inflationary potential is assumed to possess a $\phi^{2}$ bottom with $m_{\phi}=10^{-5}M_{P}$. In the left panel we set $p=0$, $c=1$, and $C_{\Upsilon}=0.1(10)$ for the solid(dashed) curves. In the right panel we set $p=0$, c=1c=-1, and $C_{\Upsilon}=2\times 10^{-8}(2\times 10^{-7})$ for the solid(dashed) curves. The end of WI is fixed at $N=0$.
@@ III Gravitational dark matter production
$+\frac{1}{M_{P}}\left[h_{\mu\nu}T^{\mu\nu}_{M}+\mathcal{L}_{3}(h)\right]+$$\Big\}.$ (14)
Here, $\tilde{\mathcal{L}}_{M}$ refers to the material fields in the FLRW background, $\mathcal{L}_{n}(h)$ refers to the $n$-th order term of $h_{\mu\nu}$, and $T^{\mu\nu}_{M}$ is the energy-momentum tensor of material fields. If we consider the perturbative quantization of gravity, then $h_{\mu\nu}$ can be regarded as graviton, and $h_{\mu\nu}T^{\mu\nu}_{M}$ describes the leading order interaction between graviton and matter.
The action Eq. (III) actually divides DM production into two channels. One is the CGPP mechanism mentioned earlier, which is induced by the DM part in $\tilde{\mathcal{L}}_{M}$. The other is the freeze-in mechanism through the graviton-mediated interactions between DM and other particles, induced by $h_{\mu\nu}T^{\mu\nu}_{M}$ terms. The interaction here is mainly the $s$-channel annihilation process, as shown in Fig. 2. During WI, both radiation particles and inflatons possess considerable energy due to the high temperature, which significantly enhances the gravitational scattering cross-section, thereby facilitating the production of massive DM particles.
@@ III.1 Cross-section of graviton-mediated annihilation
here the factor of $1/2$ is considered for neutrinos since they are Weyl particles.
@@ IV.2 Annihilation of inflatons
In contrast to the annihilation of inflatons in the SI case with the same $m_{\phi}$, the WI case exhibit two significant differences. Firstly, the DM with $m_{\chi}>m_{\phi}$ can be produced in the WI case due to the quasi-thermal state of inflatons, while it is prohibited in the SI case. Secondly, if $\phi_{e}\sim 0.1-1M_{P}$ is estimated at the end of WI, we will find that the yield of DM with $m_{\chi}\ll m_{\phi}$ in the WI case is roughly $10^{-6}$ lower than that in the SI case with $T_{\mathrm{reh}}=T_{e}$. It reveals the enormously higher annihilation efficiency of static inflatons than moving inflatons.
To understand this, we take $\rho_{r}\simeq\rho_{\phi}$ at the end of WI and rewrite Eq. (43) as:
@@
Then comparing it with Eq. (41), we observe that the interaction term for WI is suppressed by a factor of $1/g_{*}^{2}\sim 10^{-4}$. The suppression may be attributed to the reduced number density of inflatons in WI. Due to the quasi-thermal state of inflatons in WI, the number density can be estimated as $n_{\phi}\propto T^{3}\sim\rho_{\phi}/T$ with $T\gg m_{\phi}$. Conversely, in the SI scenario, it is $n_{\phi}=\rho_{\phi}/m_{\phi}$. Therefore, for the same energy density $\rho_{\phi}$, WI results in a significantly reduced inflaton number density compared to SI, thereby causing a diminished annihilation rate. Actually, $\rho_{\phi}$ at the end of WI typically takes on a lower value than that in SI, given that it usually holds that $\phi_{e}^{\mathrm{WI}}<\phi_{e}^{\mathrm{SI}}$ for the same inflationary potential. As a result, the number density $n_{\phi}$ in WI and the interaction term $R_{\phi\rightarrow\chi}^{\mathrm{WI}}$ are further suppressed by several orders of magnitude, so does the DM yield.
@@ IV.3 Contribution of CGPP mechanism
We adopt a $\phi^{2}$ WI model with mass parameters $m_{\phi}=10^{-5}M_{P}$ and $10^{-7}M_{P}$. By adjusting the parameters in Eq. (10), we obtain WI backgrounds with $T_{e}=5\times 10^{-4}M_{P}$, $H_{e}=10^{-6}M_{P}$ and $T_{e}=5\times 10^{-5}M_{P}$, $H_{e}=10^{-8}M_{P}$, respectively. The CGPP of DM with minimal coupling ($\xi=0$) and conformal coupling ($\xi=1/6$) is numerically investigated. The abundance-mass correlations are shown in Fig. 5. These results are actually indicative of the general characteristics of a wide range of WI models, although they are based on a specific model.
@@
Here we present a threshold for $T_{e}$. If WI takes place with $T_{e}$ below this threshold, there will not be an overproduction of DM for any mass. By comparison, we calculate the DM relic abundance in the $\phi^{2}$ SI model with $m_{\phi}=10^{-5}M_{P}$ and $T_{\mathrm{reh}}=3\times 10^{-5}M_{P}$, shown as the gray dashed curves in Fig. 5. It has a correlation of
$\rho_{\chi,0}\propto m_{\chi}^{0},~~~H_{\mathrm{reh}}\lesssim m_{\chi}\lesssim H_{e},$ (55)
$\rho_{\chi,0}\propto m_{\chi}^{1/2},~~~~~m_{\chi}\ll H_{\mathrm{reh}},$ (56)
for minimal coupling, and
@@
for conformal coupling, where $H_{\mathrm{reh}}$ is the Hubble parameter when reheating ends. To a certain extent, WI is similar to the SI with instantaneous reheating, corresponding to $H_{\mathrm{reh}}=H_{e}$.
Figure 5: Relic abundance of scalar DM with minimal (red curves) and conformal (blue curves) coupling via CGPP mechanism in WI. The inflationary background is served by a $\phi^{2}$ WI model. We set $m_{\phi}=10^{-5}(10^{-7})M_{P}$, $T_{e}=5\times 10^{-4}(5\times 10^{-5})M_{P}$, and $H_{e}=10^{-6}(10^{-8})M_{P}$ for the solid(dashed) curves. For comparison, the gray dashed curves depict DM relic abundance in the $\phi^{2}$ SI scenario with $m_{\phi}=10^{-5}M_{P}$ and $T_{\mathrm{reh}}=3\times 10^{-5}M_{P}$.
For the $m_{\chi}>H_{e}$ case, apparently, a little increase in mass leads to a rapid decrease in the relic abundance. The super-Hubble-mass DM with $m_{\chi}\sim 10-30H_{e}$ can be sufficiently produced and constitutes the entirety of observed DM abundance. This is similar to the SI case. Nevertheless, for the same inflationary potential, WI generally has lower $\phi_{e}$ or $H_{e}$ than SI. It actually narrows the DM mass range of overproduction compared with SI, thereby expanding the feasible DM mass range. This is clearly demonstrated in Fig. 5.
@@ V Observational constraints
where $N_{*}\sim 60$ denotes the $e$-folding number between the $k_{*}$ crossing and the end of inflation. $N_{\mathrm{tot}}$ denotes the total $e$-folding number during the inflationary phase, and its value is dependent on the choice of the infrared cut-off of $k$. Consistent with previous discussions, we set the cut-off as $k>a_{0}H_{0}$, which leads to $N_{\mathrm{tot}}\sim 65$. Combining the observational constraint and the above equations, we restrict that $\alpha>0.089$. For scalar DM with conformal coupling, this condition is inherently fulfilled, thereby eliminating any constraints on its mass. However, for the minimally coupled scalar DM, this condition implies that $m_{\chi}>0.52H_{*}\sim m_{\phi}$, which means that only DM with a super-Hubble scale mass is allowed in this case. This constraint is depicted by the red dashed line in Fig. 6. The mass ranges of DM with sufficient density are limited to $10^{-8}-10^{-2}M_{P}$ for minimal coupling and $10^{-14}-10^{-2}M_{P}$ for conformal coupling.
@@ VI Conclusion
We have investigated the purely gravitational DM production within the WI background. Three possible gravitational channels for DM production are explored, and exhibit distinctive features compared to the SI scenario. For the freeze-in channel via the SM particles annihilation, we derive the interaction term in the Boltzmann equation and calculate the DM yield. The analytical approach is utilized for the light DM case, which is consistent with the numerical results. For the inflaton annihilation channel, we obtain markedly different results from the SI scenario. Due to the quasi-thermal state of inflaton, the channel of producing super-inflaton-mass DM is open in WI, while the sub-inflaton-mass case is suppressed compared to the SI scenario. In the context of the CGPP channel, we identify the temperature threshold for sufficient DM production. We also show a correlation of $\rho_{\chi,0}\propto T_{e}^{4}m_{\chi}^{1/2}$ for minimally coupled DM and $\rho_{\chi,0}\propto m_{\chi}^{5/2}$ for conformally coupled DM, indicating differences from the SI scenario. Furthermore, the constraints on the DM mass $m_{\chi}$ and the temperature $T_{e}$ at the end of WI are explored. We demonstrate the feasible parameter space limited by the observations of DM abundance and isocurvature perturbations. The results of this study can be extended to purely gravitational fermion and vector DM, as they are similar to the scalar DM with conformal and minimal coupling, respectively.
The results of this paper may help distinguish between WI and SI if combined with the future observations of primordial gravitational waves. The energy scale of inflation at the CMB pivot scale is given by $V_{*}=\frac{3\pi^{2}}{2}r\Delta_{R}$. If the tensor-to-scalar ratio $r$ can be measured, we can confirm $V_{*}$ and estimate the energy scale when inflation ends, $V_{e}$. In WI, this means the temperature $T_{e}$ is confirmed. And the possible mass of purely gravitational DM with sufficient abundance can be read from the mass-temperature relationship we derived (shown as Fig. 6). In SI, however, the production of DM depends not only on $V_{e}$ but also on $T_{\mathrm{reh}}$, so confirming $V_{e}$ cannot restrict the DM mass with sufficient abundance. In this way, if purely gravitational DM constitutes all of the DM, one can distinguish the WI and SI by the measurement of DM mass (assuming there is a way to measure it in the future). If the measured mass align with the mass-temperature relationship of WI, it would suggest that the inflation scenario might be WI. Otherwise, if the observed mass does not align, it is more likely to be SI. The above method is very idealized, but feasible in principle.